Complete the shear moment diagram9 kN/m E B D -2 m-2 m--2m- -4.5 m -4 m
Answer #1

FBD of AB qkN/m T T RA 4.5m RB EMB=0 4. SRA= 9x4.5 x 4.5 Re 2 = 6.75 KN Ef y=0 RB - 9x4.5-6.75 13.5kn 2 FBD E aknym II TI AmA paint V = 6.75KN point V = 6.75 9x4.5 -13.5 KN 6.75X4.5 9.84.5 2 at ci x4.5 point V= -13.5_9x2=-31-5 kn (Left] 11.25 kn [Rimo n=(6.75*8-5)-(exas 454)] - (2x4x2) + (42.7542) m=40.5 kNm At point E, -RE +(9x4) = +18 KN v= m=0 At Point 18 kn a kolm RA